COMPOSITIONAL ROOTS OF HÉNON MAPS

Gregery T. Buzzard John Erik Fornaess

Abstract. Let H denote a composition of complex Hénon maps in C^{2}. In this paper we show that the only possible compositional roots of H are also compositions of Hénon maps, and that H can have compositional roots of only finitely many distinct orders.

Typeset by AMS-TEX